Consider the following snapshot of a system:

           Allocation       Max          Available 

            ABCD         ABCD           ABCD

T0        3141           6473            2224

T1        2102           4232 

T2        2413           2533 

T3        4110           6332

T4        2221           5675

Answer the following questions using the banker’s algorithm:

 a. Illustrate that the system is in a safe state by demonstrating an order in which the threads may complete.

 b. If a request from thread T4 arrives for (2, 2, 2, 4), can the request be granted immediately?

 c. If a request from thread T2 arrives for (0, 1, 1, 0), can the request be granted immediately?

 d. If a request from thread T3 arrives for (2, 2, 1, 2), can the request be granted immediately?

0 0

> Please answer the question as soon as possible..Thank you

Nisu Wed, Aug 11, 2021 6:16 AM

Add a comment Improve this question Transcribed image text
Answer #1

a) safe succession : T2, T0, T1, T3, T4

b) T4 arrive (2, 2, 2, 4) when T4 allocate available resource 7 7 7 8 + 2 1 0 2 + 4 1 1 0 = 13 9 9 10

which can fullfill the need + request = 5 6 7 5 + 2 2 2 4 = 7 8 9 9 So request granted.

c) let allocate resource to T3, after sometime available resource = 2 2 2 4 + 4 1 1 0 = 6 3 3 4 which can fullfill the need of T2 so granted only after T3.

d) When resource allocated to T3 available resource= 4 6 3 7

Which can fullfill the need and request of T3 So request granted.